A titration was performed on a 10.00-mL sample of water taken from an acidic lake. If it took 9.56 mL of 0.0412 mol/L NaOH(aq) to neutralize the sulphuric acid in the lake water sample, calculate the concentration of the sulphuric acid. What is the pH of this lake water? Full GRASS solution.


1
Expert's answer
2021-01-15T06:48:01-0500

H2SO4(aq) + 2NaOH(aq) = Na2SO4(aq) + 2H2O(l)

1 mol H2SO4 = 2 mol NaOH

no of mol of NaOH consumed = M*V

           = 0.0412"\times" 9.56

           = 0.394 mmol

no of mol of H2SO4 reacted = "\frac{0.394}{2}" = 0.197 mmol

concentration of H2SO4 in water sample = n/v

           = "\frac{0.197}{10}"

           = 0.0197 M

[H3O+] in water sample = 2"\times" 0.0197 = 0.0394 M

pH = -log[H3O+]

  = -log0.0394

  = 1.4